# Phase Space Electronic Structure Theory: From Diatomic Lambda-Doubling to Macroscopic Einstein-de Haas

### v1

@@ Abstract
$\Lambda$-doubling of diatomic molecules is a subtle microscopic phenomenon that has long attracted the attention of experimental groups, insofar as rotation of molecular nuclei induces small energetic changes in the (degenerate) electronic state. A direct description of such a phenomenon clearly requires going beyond the Born-Oppenheimer approximation. Here we show that a phase space theory previously developed to capture electronic momentum and model vibrational circular dichroism – and which we have postulated should also describe the Einstein-de Haas effect, a macroscopic manifestation of angular momentum conservation – is also able to recover the $\Lambda$-doubling energy splitting (or $\Lambda$-splitting) of the NO molecule nearly quantitatively. The key observation is that, by parameterizing the electronic Hamiltonian in terms of both nuclear position ($\bf X$}) and nuclear momentum ($\bf P$}), a phase space method yields potential energy surfaces that explicitly include the electron-rotation coupling and correctly conserve angular momentum (which we show is essential to capture $\Lambda$-doubling). The data presented in this manuscript offers another small glimpse into the rich physics that one can learn from investigating phase space potential energy surfaces $E_{PS}(\bf X,\bf P)$ as a function of both nuclear position and momentum, all at a computational cost comparable to standard Born-Oppenheimer electronic structure calculations.
@@ 1 Introduction
Conservation of angular momentum is a fundamental principle that governs both everyday life and microscopic dynamics. As far as everyday life is concerned, angular momentum conservation leads to the Coriolis force that drives the clockwise subtropical Pacific Ocean gyretalley2011descriptive in the Northern Hemisphere while also allowing a rolling hoop to stay uprightkleppner2014introduction. Because of angular momentum conservation, whenever one part of a system changes its state of motion, the remaining degrees of freedom must respond so that the total momentum and angular momentum are preservednoether1971invariant–a concept we are all well acquainted with. On a microscopic level, the same logic applies to electrons and nuclei in molecules and solids: when the nuclei move, the electrons must adjust their states to the movement of the nuclei. Just like Foucault pendulums, electrons can experience Coriolis forces.wick1948magnetic, van1951coupling, hehl1990inertial, serebrennikov2006coriolis, geilhufe2022dynamic
Unfortunately, most modern-day electronic structure theories rely on the Born-Oppenheimer approximation and ignore the direct impact of nuclear momentum on electrons bian2023total. To go beyond this limitation, we recently introduced a phase space theory that explicitly parametrizes the electronic Hamiltonian by both nuclear position and momentum and conserves the total momentum and angular momentum of electrons and nuclei. We have already demonstrated that the method can recover vibrational circular dichroismduston2024phase, tao2024electronic and Raman optical activitytao2025non; for model systems, the approach can also offer improved vibrational energiesbian2025phase, wu2025recovering. Moreover, because angular momentum conservation is guaranteed, we have postulated that the approach should be useful for modeling chiral-induced spin selectivitybian2025review as well as the Einstein-de Haas (EdH) effect einstein1915experimenteller, bradbury2025symmetry. The latter phenomenon remains of great interest to physicists and chemists alikementink2019quantum, ganzhorn2016quantum; as a brief review, EdH is the phenomenon whereby changing the magnetic moment can cause a ferromagnet to rotate, and thus serves as a macroscopically large example of angular momentum conservation in solids.
In the realm of small molecule chemistry, here we will now show that the same beyond BO physical concepts described above also naturally describe $\Lambda$-doubling. In particular, using the phase space electronic structure theory from Refs.qiu2024simple, tao2024practical, tao2025basis, we will show that one naturally recovers the $\Lambda$-doubling splittings of the NO molecule quantitatively. Thus, we will argue that, over the smallest or largest length scales, electronic structure theory can be improved by moving to a phase space framework where the Schrodinger equation is solved in the frame of moving nuclei.
@@ 2 Brief Review of the Phase Space Method
Before we address $\Lambda$-doubling for NO, let us briefly review phase space electronic structure theory. Phase space electronic structure theory provides a semiclassical means to go beyond the standard BO approximation by parametrizing the electronic Hamiltonian using both the classical nuclear position ($\mathbf{X}$) and nuclear momentum ($\mathbf{P}$). More than ten years ago, Shenvi wrote down the first such example shenvi2009phase, suggesting that we diagonalize:
@@
is a local version of moment of inertia partitioned to each atom, and $\mathcal{I}_{3}$ is the $3\times 3$ identity matrix. For the case of a linear molecule (as below) where $\mathbf{K}_{B}$ is not invertible, we replace $\mathbf{K}_{B}^{-1}\hat{\mathbf{J}}_{B}$ with qiu2024simple:
@@
For example, when the symmetry constraints above are satisfied, the coupling term exactly captures the non-inertial Coriolis effect on electrons in the rotating frame of rigid nuclei. To verify this statement, imagine that we include $\hat{\mathbf{\Gamma}}_{A}^{\prime}$ and $\hat{\mathbf{\Gamma}}_{A}^{\prime\prime}$ in the $\hat{\mathbf{\Gamma}}_{A}$ operator (and thus satisfy a spinless version of Eq. 4, i.e., we calculate the spatial part of the electrons in the body frame while the electronic spins are in the space-fixed frame); the relevant $\hat{\mathbf{\Gamma}}_{A}$ operator satisfies $i\hbar\sum_{A}\mathbf{X}_{A}\times\hat{\mathbf{\Gamma}}_{A}+\hat{\mathbf{L}}^{e}=0$ and conserves the total angular momentum excluding spinqiu2024simple. When the nuclei are rotating around the origin (we set the center of mass as the origin) with an angular velocity $\omega$, the coupling term produced by the $\hat{\mathbf{\Gamma}}$ becomes
$-\sum_{A}\frac{i\hbar}{M_{A}}\mathbf{P}_{A}\cdot(\hat{\mathbf{\Gamma}}_{A}^{\prime}(\mathbf{X})+\hat{\mathbf{\Gamma}}_{A}^{\prime\prime}(\mathbf{X}))=$ $-\sum_{A}i\hbar\left(\mathbf{\omega}\times\mathbf{X}_{A}\right)\cdot(\hat{\mathbf{\Gamma}}_{A}^{\prime}(\mathbf{X})+\hat{\mathbf{\Gamma}}_{A}^{\prime\prime}(\mathbf{X}))$
@@
which is exactly the electronic Coriolis term in a rotating non-inertial frame of an angular velocity $\mathbf{\omega}$. Previously, this result has been shown to hold generally for any type of nuclear motion, including both rotations and vibrations, in the nonlocal limit tao2025basis, qiu2024simple (where $\sigma\rightarrow\infty$ in Eq. 6); here, we emphasize that this result holds regardless of the choice of the locality parameter in the case of pure rotations. Moreover, further including the electron spin rotation factor $\hat{\mathbf{\Gamma}}_{A}^{\prime\prime\prime}$ so as to satisfy the original Eq. 4 (corresponding to a calculation of the spatial and spin components of electrons in the body frame), we find a coupling term
@@ 3 Phase Space in the Case of Degeneracy
One of the most interesting facets of phase space electronic structure theory is the behavior of potential energy surfaces in the presence of degeneracy. In this context, a phase space approach to electronic structure can break symmetries that are protected within BO theory. In particular, with the BO framework, every radical molecule (with one unpaired electron) must be part of a degenerate electronic subspace (usually a doublet) according to Kramers’ theoremkramers1930theorie. That being said, however, according to a phase space description, the two electronic states may interact differently with nuclear motion (that breaks time reversal symmetry) and lose their energetic degeneracy. As a result, even for the smallest molecule, one finds multiple minima in $\bf P$}-space; see Ref. bradbury2025symmetry.
3.1 Multiple Minima on the Potential Energy Surfaces
@@
In Fig. 1, we plot the two lowest phase space surfaces from solving Eq. 2 for the molecule NO in a STO-3G basis. On the $x$-axis, our order parameter is the canonical nuclear rotation angular momentum $\mathbf{L}=\sum_{A}\mathbf{X}_{A}\times\mathbf{P}_{A}$. Note that, whereas we find a doubly degenerate ground state PES under the BO framework, PS theory now splits into two shifted, approximately parabolic phase space PESs. As will be discussed below in Sec. 4.1, the new broken-symmetry eigenstates have different parities in the case of diatomics. Note, however, that the two PESs are still degenerate at zero canonical nuclear momentum and have a mirror symmetry with respect to $\mathbf{L}=0$, all due to time reversal symmetry in the combined electronic and nuclear space.
In order to understand the physical meaning of the two split PESs and of their unique minima along the canonical nuclear rotational angular momentum $\mathbf{L}$, one must recognize that the variable $\mathbf{P}$ – which would appear to be the nuclear linear momentum within the Born-Oppenheimer framework – represents instead the combined nuclear and electronic momentum littlejohn2024diagonalizing. This representation can be shown by applying the total momentum operator $\hat{\mathbf{P}}+\hat{\mathbf{p}}=\sum_{A}\hat{\mathbf{P}}_{A}+\sum_{i}\hat{\mathbf{p}}_{i}$ to a coupled electronic and nuclear wavefunction $\Psi(\mathbf{X},\mathbf{\mathbf{r}})=\sum_{n}\psi_{n}(\mathbf{X})\phi_{n}(\mathbf{X},\mathbf{\mathbf{r}})$ in the BO formalism
@@
where ${\hat{\bf d}}_{A}$ is the derivative coupling operator, one can intuitively think of ${\hat{\bf d}}_{A}$ as incorporating the electronic momentumlittlejohn2023representation, littlejohn2024diagonalizing. Thus, while ${\hat{\bf P}}_{A}$ represents the total electronic + nuclear momentum, ${\hat{\bf P}}_{A}-i\hbar{\hat{\bf d}}_{A}$ represents the nuclear momentum. Analogous statements apply to angular momentum as well as linear momentum: the canonical nuclear rotational angular momentum $\mathbf{L}$ in fact represents the total angular momentum $\mathbf{J}=\mathbf{R}+\mathbf{L}^{e}+\mathbf{S}^{e}$ where $\mathbf{R}$ is the nuclear rotational angular momentum, $\mathbf{L}^{e}$ and $\mathbf{S}^{e}$ are the electronic orbital and spin angular momentum vanHorn2025electronic.
@@
3.2 Extracting Rotational Energies
Within BO theory, the simplest means of extracting vibrational and rotational energies is very well exploredbarone2011computational, wilson1980molecular, mccoy1992algebraic. Namely, one finds the local minimum geometry ${\bf X}_{0}$ of $E_{BO}({\bf X}_{0})$ in nuclear configuration space, one separates rotations from translations by looking for zeros of the Hessian, one makes a rigid rotor approximation for the rotations, and one makes a harmonic approximation for the vibrations.
@@ 3.2 Extracting Rotational Energies
Eq. 22 is obviously quite reminiscent of more advanced theories of molecular dynamics that separate rotations and vibrations; for example, see Ref. bunker2006molecular. For this article (with a focus on the spectrum of NO), we will focus exclusively on rotations; vibrations and vibrational-rotational coupling will be treated within a forthcoming article. With the interpretation above, the rotational kinetic energy of a molecular system is written as
@@
The operator evaluated in Eq. 24 is often termed the $L$-uncoupling and $S$-uncoupling operatorslefebvre2004spectra. In the context of ground state diatomics, as we will show below (Sec. 4.3), such an electron-nuclear coupling exactly gives rise to the $\Lambda$-doubling effect.
@@ 4 $\Lambda$-doubling and NO
Strictly speaking, for a diatomic molecule, as for any molecule, the good quantum numbers are $J$, the total angular momentum quantum number, and its lab-frame $z$-axis projection $M$; each energy eigenstate can be simultaneously labeled by one $J$ and one $M$. The $\Lambda$-doubling effect occurs when there are two degenerate electronic configurations with positive and negative angular momentum projections $\Lambda$ along the internuclear axis according to the BO electronic Hamiltonian. As will be seen below, this degeneracy is lost due to electron-rotation coupling, leading to a splitting in the rotational spectrum.
4.1 Standard View of $\Lambda$-doubling
@@
where $D$ is the Wigner $D$-matrix Varshalovich1988. Note that $J$ and $M$ are the only good quantum numbers for an isolated system that has rotational invariance. Additionally included in Eq. 4.1 is an energy label $n$ ($n=X,A,B$...) and a separable vibrational wavefunction $|\nu\rangle$, but these parts will be omitted in the following to simplify the discussion.
@@
Here, $p^{\pm}$ refers to the parity eigenstate with total parity $\pm(-1)^{J-S}$ for electronic $\Pi$ terms. For the half-integer J values and $S=\frac{1}{2}$ considered here, the $|p^{+}\rangle$ state with parity $(-1)^{J-\frac{1}{2}}$ is further labeled $e$, while the $|p^{-}\rangle$ state with parity $(-1)^{J-\frac{1}{2}}$ is labeled $f$. $A^{\prime}$ and $A^{\prime\prime}$ have also been used for labeling based on the reflection symmetry of the electronic spatial coordinates. alexander1988nomenclature
Now, if we consider the effect of nuclear rotations as in Eq. 25, the $^{2}\Pi$ manifold interacts with higher-lying $^{2}\Sigma$ states. (Importantly, the $\Sigma$ label here signifies $\Lambda=0$; in an unfortunate aspect of small molecule notation, $\Sigma$ has two different meanings and the $\Sigma$ here is different from the $\Sigma$ in Eq. 4.1 which refers to the internuclear axis projection of the spin angular momentum.) For NO, the interaction with high-lying $^{2}\Sigma$ states involves the off-diagonal $L$-uncoupling operatorlefebvre2004spectra,
@@
Evaluating Eq. 31 is difficult as one needs to include enough states for convergence (which is many). Finally, note that the $S$-uncoupling operator in Eq. 24 also couples together the $^{2}\Pi_{1/2}$ and $^{2}\Pi_{3/2}$ manifolds for $J\geq 3/2$. A careful diagonalization within the combined space of $^{2}\Pi_{1/2}$ and $^{2}\Pi_{3/2}$ manifolds zare1988angular, hinkley1972lambda yields the following result for the $\Lambda$-splitting of the $^{2}\Pi_{1/2}$ manifold
@@
This concludes our brief review of the standard, well-established formal theory of $\Lambda$-doublingbrown2003rotational, zare1988angular, bunker2006molecular. Note that below we will compare phase space theory directly to experimental data (rather than to Eqs. 32-33), but the analysis above will be helpful for reasons of interpretation. Note also that the derivation of Eq. 32 ignores the parity dependency of the energy denominator, which would require a much more involved analysis to include gordon2025energy.
4.2 A 1D Model Hamiltonian Compatible with Phase Space Electronic Structure Theory
Eqs. 31-33 above represent the standard, perturbative approach to $\Lambda$-doubling that one finds in the literature. We would now like to show that this result can also be recapitulated quite naturally (and without a sum over states) using PS theory; moreover, the PS approach is not based on perturbation theory.
In line with Sec. 3.2 above, the simplest model Hamiltonian that can capture coupled nuclear rotation-electronic effects is a standard rigid rotor that can be rotated with angle $\theta$ around one axis (hereby the name “1D” model). We let $\theta$ be the rotational angle of interest and we imagine that there are two electronic states of interest (as in the case of NO described above in Fig. 1) with $\pm\alpha$ total electronic angular momentum. In the spirit of Eq. 20, the kinetic momentum is $\frac{\hbar}{i}\frac{\partial}{\partial\theta}\pm\alpha$, which leads to the following (trivially separable) Schrodinger equation:
@@ 4.2 A 1D Model Hamiltonian Compatible with Phase Space Electronic Structure Theory
Within a perturbative treatment, $|\Psi_{\mathrm{SOC}}^{(1)}\rangle$ should be a parity eigenstate. To a good approximation, we can also write Eq. 4.2 above as
@@
4.3 A 2D Model Hamiltonian Compatible with Phase Space Electronic Structure Theory
Next, let us imagine using phase space electronic structure theory to describe $\Lambda$-doubling in a full three dimensional context, beyond the 1D model in Sec. 4.2 above (which is trivially solvable). As we will see, a multidimensional model cannot be solved analytically and requires either a large matrix diagonalization or approximations.
To begin our multidimensionl analysis, let $c$ denote the molecular axis of the diatomic molecule (from N to O). Due to the $C_{\infty v}$ symmetry of the NO molecule around the internuclear $c$-axis, any choice of an axis perpendicular to $c$ is equivalent. We will pick the $a$- and $b$-axes associated with $\chi=0$ zare1988angular, zare1973direct, where $\chi$ is the last of the Euler angles $(\phi,\theta,\chi)$ that describes the molecular orientation in the zzyyzz convention. As can be found in many textbooks bunker2006molecular, brown2003rotational, the quantum mechanical Hamiltonian for the nuclear kinetic energy associated with rotations of a linear molecule
@@ 4.3 A 2D Model Hamiltonian Compatible with Phase Space Electronic Structure Theory
The Hamiltonian $\hat{H}_{\mathrm{rot}}^{\mathrm{iso}}$ has one more degree of freedom than the true Hamiltonian in Eq. 43; namely $\chi$ remains as an independent variable in the definition of $\hat{\mathbf{J}}$. For an isomophormism to hold, we must diagonalize Eq. 44 within a basis that imposes the condition that $J_{c}=J_{c}^{e}$, so as to remove all extraneous solutions; for more details, see Refs. brown2003rotational, zare1988angular.
@@
and $|\hat{J}$$|$ denotes the 2-norm (square root of the sum of squares) of operator $\hat{J}$ . At this point, we can rewrite the coupling term as a matrix

### v2

@@ Abstract
$\Lambda$-doubling of diatomic molecules is a subtle microscopic phenomenon that has long attracted the attention of experimental groups, insofar as rotation of molecular nuclei induces small energetic changes in the (degenerate) electronic state. A direct description of such a phenomenon clearly requires going beyond the Born-Oppenheimer approximation. Here we show that a phase space theory previously developed to capture electronic momentum and model vibrational circular dichroism – and which we have postulated should also describe the Einstein-de Haas effect, a macroscopic manifestation of angular momentum conservation – is also able to recover the $\Lambda$-doubling energy splitting (or $\Lambda$-splitting) of the NO molecule nearly quantitatively and nonperturbatively (without a sum over states). The key observation is that, by parameterizing the electronic Hamiltonian in terms of both nuclear position ($\bf X$}) and nuclear momentum ($\bf P$}), a phase space method yields potential energy surfaces that explicitly include the electron-rotation coupling and correctly conserve angular momentum (which we show is essential to capture $\Lambda$-doubling). The data presented in this manuscript offer another small glimpse into the rich physics that one can learn from investigating phase space potential energy surfaces $E_{PS}(\bf X,\bf P)$ as a function of both nuclear position and momentum, all at a computational cost comparable to standard Born-Oppenheimer electronic structure calculations.
@@ 1 Introduction
Conservation of angular momentum is a fundamental principle that governs both everyday life and microscopic dynamics. As far as everyday life is concerned, angular momentum conservation leads to the Coriolis force that drives the clockwise subtropical Pacific Ocean gyre58 in the Northern Hemisphere while also allowing a rolling hoop to stay upright34. Because of angular momentum conservation, whenever one part of a system changes its state of motion, the remaining degrees of freedom must respond so that the total momentum and angular momentum are preserved47–a concept we are all well acquainted with. On a microscopic level, the same logic applies to electrons and nuclei in molecules and solids: when the nuclei move, the electrons must adjust their states to the movement of the nuclei. Just like Foucault pendulums, electrons can experience Coriolis forces.66, 64, 27, 51, 24
Unfortunately, most modern-day electronic structure theories rely on the Born-Oppenheimer approximation and ignore the direct impact of nuclear momentum on electrons 8. To go beyond this limitation, we recently introduced a phase space theory that explicitly parametrizes the electronic Hamiltonian by both nuclear position and momentum and conserves the total momentum and angular momentum of electrons and nuclei. We have already demonstrated that the method can recover vibrational circular dichroism20, 60 and Raman optical activity59; for model systems, the approach can also offer improved vibrational energies7, 68. Moreover, because angular momentum conservation is guaranteed, we have postulated that the approach should be useful for modeling chiral-induced spin selectivity6 as well as the Einstein-de Haas (EdH) effect 21, 11. The latter phenomenon remains of great interest to physicists and chemists alike44, 23; as a brief review, EdH is the phenomenon whereby changing the magnetic moment can cause a ferromagnet to rotate, and thus serves as a macroscopically large example of angular momentum conservation in solids.
In the realm of small molecule chemistry, here we will now show that the same beyond BO physical concepts described above also naturally describe $\Lambda$-doubling. In particular, using the phase space electronic structure theory from Refs.49, 61, 62, we will show that one naturally recovers the $\Lambda$-doubling splittings of the NO molecule quantitatively. Thus, we will argue that, over the smallest or largest length scales, electronic structure theory can be improved by moving to a phase space framework where the Schrodinger equation is solved in the frame of moving nuclei.
@@ 2 Brief Review of the Phase Space Method
Before we address $\Lambda$-doubling for NO, let us briefly review phase space electronic structure theory. Phase space electronic structure theory provides a semiclassical means to go beyond the standard BO approximation by parametrizing the electronic Hamiltonian using both the classical nuclear position ($\mathbf{X}$) and nuclear momentum ($\mathbf{P}$). More than ten years ago, Shenvi wrote down the first such example 52, suggesting that we diagonalize:
@@
is a local version of moment of inertia partitioned to each atom, and $\mathcal{I}_{3}$ is the $3\times 3$ identity matrix. For the case of a linear molecule (as below) where $\mathbf{K}_{B}$ is not invertible, we replace $\mathbf{K}_{B}^{-1}\hat{\mathbf{J}}_{B}$ with 49:
@@
For example, when the symmetry constraints above are satisfied, the coupling term exactly captures the non-inertial Coriolis effect on electrons in the rotating frame of rigid nuclei. To verify this statement, imagine that we include $\hat{\mathbf{\Gamma}}_{A}^{\prime}$ and $\hat{\mathbf{\Gamma}}_{A}^{\prime\prime}$ in the $\hat{\mathbf{\Gamma}}_{A}$ operator (and thus satisfy a spinless version of Eq. 4, i.e., we calculate the spatial part of the electrons in the body frame while the electronic spins are in the space-fixed frame); the relevant $\hat{\mathbf{\Gamma}}_{A}$ operator satisfies $i\hbar\sum_{A}\mathbf{X}_{A}\times\hat{\mathbf{\Gamma}}_{A}+\hat{\mathbf{L}}^{e}=0$ and conserves the total angular momentum excluding spin49. When the nuclei are rotating around the origin (we set the center of mass as the origin) with an angular velocity $\omega$, the coupling term produced by the $\hat{\mathbf{\Gamma}}$ becomes
$-\sum_{A}\frac{i\hbar}{M_{A}}\mathbf{P}_{A}\cdot(\hat{\mathbf{\Gamma}}_{A}^{\prime}(\mathbf{X})+\hat{\mathbf{\Gamma}}_{A}^{\prime\prime}(\mathbf{X}))$
$=$ $-\sum_{A}i\hbar\left(\mathbf{\omega}\times\mathbf{X}_{A}\right)\cdot(\hat{\mathbf{\Gamma}}_{A}^{\prime}(\mathbf{X})+\hat{\mathbf{\Gamma}}_{A}^{\prime\prime}(\mathbf{X}))$
@@
which is exactly the electronic Coriolis term in a rotating non-inertial frame of an angular velocity $\mathbf{\omega}$. Previously, this result has been shown to hold generally for any type of nuclear motion, including both rotations and vibrations, in the nonlocal limit 62, 49 (where $\sigma\rightarrow\infty$ in Eq. 6); here, we emphasize that this result holds regardless of the choice of the locality parameter in the case of pure rotations. Moreover, further including the electron spin rotation factor $\hat{\mathbf{\Gamma}}_{A}^{\prime\prime\prime}$ so as to satisfy the original Eq. 4 (corresponding to a calculation of the spatial and spin components of electrons in the body frame), we find a coupling term
@@ 3 Phase Space in the Case of Degeneracy
One of the most interesting facets of phase space electronic structure theory is the behavior of potential energy surfaces in the presence of degeneracy. In this context, a phase space approach to electronic structure can break symmetries that are protected within BO theory. In particular, with the BO framework, every radical molecule (with one unpaired electron) must be part of a degenerate electronic subspace (usually a doublet) according to Kramers’ theorem35. That being said, however, according to a phase space description, the two electronic states may interact differently with nuclear motion (that breaks time reversal symmetry) and lose their energetic degeneracy. As a result, even for the smallest molecule, one finds multiple minima in $\bf P$}-space; see Ref. 11.
Multiple Minima on the Potential Energy Surfaces
@@
In Fig. 1, we plot the two lowest phase space surfaces from solving Eq. 2 for the molecule NO in a STO-3G basis. On the $x$-axis, our order parameter is the canonical nuclear rotation angular momentum $\mathbf{L}=\sum_{A}\mathbf{X}_{A}\times\mathbf{P}_{A}$. Note that, whereas we find a doubly degenerate ground state PES under the BO framework, PS theory now splits into two shifted, approximately parabolic phase space PESs. As will be discussed below in Sec. 4, the new broken-symmetry eigenstates have different parities in the case of diatomics. Note, however, that the two PESs are still degenerate at zero canonical nuclear momentum and have a mirror symmetry with respect to $\mathbf{L}=0$, all due to time reversal symmetry in the combined electronic and nuclear space.
In order to understand the physical meaning of the two split PESs and of their unique minima along the canonical nuclear rotational angular momentum $\mathbf{L}$, one must recognize that the variable $\mathbf{P}$ – which would appear to be the nuclear linear momentum within the Born-Oppenheimer framework – represents instead the combined nuclear and electronic momentum 40. This representation can be shown by applying the total momentum operator $\hat{\mathbf{P}}+\hat{\mathbf{p}}=\sum_{A}\hat{\mathbf{P}}_{A}+\sum_{i}\hat{\mathbf{p}}_{i}$ to a coupled electronic and nuclear wavefunction $\Psi(\mathbf{X},\mathbf{\mathbf{r}})=\sum_{n}\psi_{n}(\mathbf{X})\phi_{n}(\mathbf{X},\mathbf{\mathbf{r}})$ in the BO formalism
@@
where ${\hat{\bf d}}_{A}$ is the derivative coupling operator, one can intuitively think of ${\hat{\bf d}}_{A}$ as incorporating the electronic momentum39, 40. Thus, while ${\hat{\bf P}}_{A}$ represents the total electronic + nuclear momentum, ${\hat{\bf P}}_{A}-i\hbar{\hat{\bf d}}_{A}$ represents the nuclear momentum. Analogous statements apply to angular momentum as well as linear momentum: the canonical nuclear rotational angular momentum $\mathbf{L}$ in fact represents the total angular momentum $\mathbf{J}=\mathbf{R}+\mathbf{L}^{e}+\mathbf{S}^{e}$ where $\mathbf{R}$ is the nuclear rotational angular momentum, $\mathbf{L}^{e}$ and $\mathbf{S}^{e}$ are the electronic orbital and spin angular momentum 63.
@@
Extracting Rotational Energies
Within BO theory, the simplest means of extracting vibrational and rotational energies is very well explored3, 67, 43. Namely, one finds the local minimum geometry ${\bf X}_{0}$ of $E_{BO}({\bf X}_{0})$ in nuclear configuration space, one separates rotations from translations by looking for zeros of the Hessian, one makes a rigid rotor approximation for the rotations, and one makes a harmonic approximation for the vibrations.
@@
Eq. 22 is obviously quite reminiscent of more advanced theories of molecular dynamics that separate rotations and vibrations; for example, see Ref. 13. For this article (with a focus on the spectrum of NO), we will focus exclusively on rotations; vibrations and vibrational-rotational coupling will be treated within a forthcoming article. With the interpretation above, the rotational kinetic energy of a molecular system is written as
@@
The operator evaluated in Eq. 24 is often termed the $L$-uncoupling and $S$-uncoupling operators37. In the context of ground state diatomics, as we will show below (Sec. 4), such an electron-nuclear coupling exactly gives rise to the $\Lambda$-doubling effect.
@@ 4 $\Lambda$-doubling and NO
Strictly speaking (i.e. without the Born-Oppenheimer approximation), for a diatomic molecule, as for any molecule, the only good quantum numbers are $J$, the total angular momentum quantum number, and its lab-frame $z$-axis projection $M$; each energy eigenstate can be simultaneously labeled by one $J$ and one $M$. The $\Lambda$-doubling effect occurs when there are two degenerate electronic configurations with positive and negative angular momentum projections $\Lambda$ along the internuclear axis according to the BO electronic Hamiltonian. As will be seen below, this degeneracy is lost due to electron-rotation coupling, leading to a splitting in the rotational spectrum.
Standard View of $\Lambda$-doubling
@@
where $D$ is the Wigner $D$-matrix 65. Note that $J$ and $M$ are the only good quantum numbers for an isolated system that has rotational invariance. Additionally included in Eq. 4 is an energy label $n$ ($n=X,A,B$...) and a separable vibrational wavefunction $|\nu\rangle$, but these parts will be omitted in the following to simplify the discussion.
@@
Here, $p^{\pm}$ refers to the parity eigenstate with total parity $\pm(-1)^{J-S}$ for electronic $\Pi$ terms. For the half-integer J values and $S=\frac{1}{2}$ considered here, the $|p^{+}\rangle$ state with parity $(-1)^{J-\frac{1}{2}}$ is further labeled $e$, while the $|p^{-}\rangle$ state with parity $(-1)^{J-\frac{1}{2}}$ is labeled $f$. $A^{\prime}$ and $A^{\prime\prime}$ have also been used for labeling based on the reflection symmetry of the electronic spatial coordinates. 1
Now, if we consider the effect of nuclear rotations as in Eq. 25, the $^{2}\Pi$ manifold interacts with higher-lying $^{2}\Sigma$ states. (Importantly, the $\Sigma$ label here signifies $\Lambda=0$; in an unfortunate aspect of small molecule notation, $\Sigma$ has two different meanings and the $\Sigma$ here is different from the $\Sigma$ in Eq. 4 which refers to the internuclear axis projection of the spin angular momentum.) For NO, the interaction with high-lying $^{2}\Sigma$ states involves the off-diagonal $L$-uncoupling operator37,
@@
Evaluating Eq. 31 is difficult as one needs to include enough states for convergence (which is many). Finally, note that the $S$-uncoupling operator in Eq. 24 also couples together the $^{2}\Pi_{1/2}$ and $^{2}\Pi_{3/2}$ manifolds for $J\geq 3/2$. A careful diagonalization within the combined space of $^{2}\Pi_{1/2}$ and $^{2}\Pi_{3/2}$ manifolds 70, 28 yields the following result for the $\Lambda$-splitting of the $^{2}\Pi_{1/2}$ manifold
@@
This concludes our brief review of the standard, well-established formal theory of $\Lambda$-doubling12, 70, 13. Note that below we will compare phase space theory directly to experimental data (rather than to Eqs. 32-33), but the analysis above will be helpful for reasons of interpretation. Note also that the derivation of Eq. 32 ignores the parity dependency of the energy denominator, which would require a much more involved analysis to include 26.
A 1D Model Hamiltonian Compatible with Phase Space Electronic Structure Theory
Eqs. 31-33 above represent the standard, perturbative approach to $\Lambda$-doubling that one finds in the literature. We would now like to show that this result can also be recapitulated quite naturally (and without a sum over states) using PS theory; the PS approach is not based on perturbation theory.
In line with Sec. 3 above, the simplest model Hamiltonian that can capture coupled nuclear rotation-electronic effects is a standard rigid rotor that can be rotated with angle $\theta$ around one axis (hereby the name “1D” model). We let $\theta$ be the rotational angle of interest and we imagine that there are two electronic states of interest (as in the case of NO described above in Fig. 1) with $\pm\alpha$ total electronic angular momentum. In the spirit of Eq. 20, the kinetic momentum is $\frac{\hbar}{i}\frac{\partial}{\partial\theta}\pm\alpha$, which leads to the following (trivially separable) Schrodinger equation:
@@
Within a perturbative treatment, $|\Psi_{\mathrm{SOC}}^{(1)}\rangle$ should be a parity eigenstate. To a good approximation, we can also write Eq. 4 above as
@@
A 2D Model Hamiltonian Compatible with Phase Space Electronic Structure Theory
Next, let us imagine using phase space electronic structure theory to describe $\Lambda$-doubling in a full three dimensional context, beyond the 1D model above (which is trivially solvable). As we will see, a multidimensional model cannot be solved analytically and requires either a large matrix diagonalization or approximations.
To begin our multidimensionl analysis, let $c$ denote the molecular axis of the diatomic molecule (from N to O). Due to the $C_{\infty v}$ symmetry of the NO molecule around the internuclear $c$-axis, any choice of an axis perpendicular to $c$ is equivalent. We will pick the $a$- and $b$-axes associated with $\chi=0$ 70, 71, where $\chi$ is the last of the Euler angles $(\phi,\theta,\chi)$ that describes the molecular orientation in the zzyyzz convention. As can be found in many textbooks 13, 12, the quantum mechanical Hamiltonian for the nuclear kinetic energy associated with rotations of a linear molecule
@@
The Hamiltonian $\hat{H}_{\mathrm{rot}}^{\mathrm{iso}}$ has one more degree of freedom than the true Hamiltonian in Eq. 43; namely $\chi$ remains as an independent variable in the definition of $\hat{\mathbf{J}}$. For an isomophormism to hold, we must diagonalize Eq. 44 within a basis that imposes the condition that $J_{c}=J_{c}^{e}$, so as to remove all extraneous solutions; for more details, see Refs. 12, 70.
@@
and $|\hat{J}$$|$ denotes the 2-norm (square root of the sum of squares) of operator $\hat{J}$. At this point, we can rewrite the coupling term as a matrix in the bases of $|\Lambda=1,S=1/2,\Sigma=-1/2\rangle$ and $|\Lambda=-1,S=1/2,\Sigma=1/2\rangle$ electronic states. Because we find the matrices of $\hat{J}^{e}_{a}$ and of $\hat{J}^{e}_{b}$ are real off-diagonal and complex off-diagonal in such bases with a certain phase choice, the coupling term becomes